# What Matters When Repurposing Diffusion Models for General Dense Perception Tasks?

### v3

@@ Appendix B More Visualization Analysis
Quantitative Comparisons of Ablations. Visualization of the ablation study experiments in the main paper §2 is shown in fig. 7. The estimated depth detail remains comparable with a customized “DPT head”. Models trained with lower-quality data like “Taskonomy + Cityscapes” or without the pre-trained VAE decoder parameters suffer from a decline in the ability to predict details. We attribute them to the low-quality annotation and the oversized decoder, respectively.
@@
Forward Diffusion Process. More detailed visualization of different noise forms and proportions in the forward diffusion process is shown in fig. 8. With larger $\beta$ values of the DDIM scheduler, the proportion of noise will be much higher during the forward diffusion process. When $\beta_{start},\beta_{end}$ reaches (1.0, 1.0), the noisy latent will be pure noise latent, which is Gaussian noise and RGB latent for (a) and (b), respectively.
Appendix C An Extra Attempt on Human Pose Estimation
@@
Appendix D Implementation Details
Implementation Details of Preliminary Section in the Main Paper. Unless specified otherwise, we train for 30000 iterations to estimate the affine-invariant monocular depth on a mixture of the Hypersim (Roberts et al., 2021) (90%) and Virtual KITTI (Cabon et al., 2020) (10%) datasets. We freeze the VAE AutoEncoder and fine-tune the U-Net of Stable Diffusion v2.1 to estimate the ground-truth label latent, with a resolution of (480, 640), a batch size of 32, and a learning rate of 3e-5. The multi-resolution noise (Ke et al., 2024) is employed for Gaussian noise and not used for RGB noise (Lee et al., 2024) by default. For inference, the ensemble size and denoising steps are set to 1 and 10, respectively. Evaluation results of absolute relative error (AbsRel) and $\delta_{1}$ are reported on five unseen monocular depth datasets, including KITTI (Geiger et al., 2013), NYU (Silberman et al., 2012), ScanNet (Dai et al., 2017), DIODE (Vasiljevic et al., 2019), and ETH3D (Schops et al., 2017).
Appendix E More Quantitative and Qualitive Experiments
E.1 Monocular Depth Estimation
@@
E.2 Surface Normal Estimation
@@
E.3 Dichotomous Image Segmentation
More Quantitative Evaluation. To conduct a comprehensive evaluation, we compare our approach with numerous previous models including models for medical image segmentation (Ronneberger et al., 2015), salient object detection (Qin et al., 2019; Zhao et al., 2020; Wei et al., 2020; Chen et al., 2020; Qin et al., 2020), camouflaged object detection (Fan et al., 2021a; Mei et al., 2021), semantic segmentation (Zhao et al., 2017; Chen et al., 2018a; Wang et al., 2020; Yu et al., 2018; Zhao et al., 2018; Howard et al., 2019; Fan et al., 2021c; Nirkin et al., 2021) and the model IS-Net (Qin et al., 2022) specifically trained for DIS. As shown in table 12, our proposed model significantly outperforms other methods across most evaluation metrics on this challenging dataset. The results highlight the effectiveness of our approach for DIS.
@@ E.3 Dichotomous Image Segmentation
DIS-VD $maxF_{\beta}\uparrow$ 0.692 0.731 0.678 0.685 0.648 0.748 0.665 0.691 0.691 0.660 0.726 0.662 0.697 0.714 0.696 0.734 0.791 0.857
$F^{w}_{\beta}\uparrow$ 0.586 0.641 0.574 0.595 0.542 0.656 0.584 0.604 0.603 0.568 0.641 0.548 0.609 0.642 0.613 0.640 0.717 0.835
${}M~{}\downarrow$ 0.113 0.094 0.110 0.107 0.118 0.090 0.110 0.106 0.102 0.114 0.095 0.116 0.102 0.092 0.103 0.096 0.074 0.040
$S_{\alpha}\uparrow$ 0.745 0.768 0.723 0.733 0.718 0.781 0.727 0.740 0.744 0.716 0.767 0.728 0.747 0.758 0.740 0.773 0.813 0.870
$E_{\phi}^{m}\uparrow$ 0.785 0.816 0.783 0.800 0.765 0.823 0.798 0.811 0.802 0.796 0.824 0.767 0.811 0.841 0.817 0.814 0.856 0.934
${\rm HCE}_{\gamma}\downarrow$ 1337 1402 1493 1567 1555 1413 1568 1606 1588 1520 1560 1660 1503 1625 1598 1324 1116 1511
DIS-TE1 $maxF_{\beta}\uparrow$ 0.625 0.688 0.620 0.640 0.598 0.694 0.644 0.646 0.645 0.601 0.668 0.595 0.631 0.669 0.648 0.695 0.740 0.841
$F^{w}_{\beta}\uparrow$ 0.514 0.595 0.517 0.549 0.495 0.601 0.558 0.552 0.557 0.506 0.579 0.474 0.535 0.595 0.562 0.597 0.662 0.814
${}M~{}\downarrow$ 0.106 0.084 0.099 0.095 0.103 0.083 0.094 0.094 0.089 0.102 0.088 0.108 0.095 0.083 0.090 0.082 0.074 0.038
$S_{\alpha}\uparrow$ 0.716 0.754 0.701 0.721 0.705 0.760 0.727 0.722 0.725 0.694 0.742 0.695 0.716 0.740 0.723 0.761 0.787 0.868
$E_{\phi}^{m}\uparrow$ 0.750 0.801 0.766 0.783 0.750 0.801 0.791 0.786 0.791 0.772 0.797 0.741 0.784 0.818 0.798 0.803 0.820 0.918
${\rm HCE}_{\gamma}\downarrow$ 233 220 230 244 271 224 274 253 267 234 262 288 234 274 249 205 149 204
DIS-TE2 $maxF_{\beta}\uparrow$ 0.703 0.755 0.702 0.712 0.673 0.756 0.700 0.720 0.724 0.681 0.747 0.680 0.716 0.743 0.720 0.759 0.799 0.876
$F^{w}_{\beta}\uparrow$ 0.597 0.668 0.598 0.620 0.570 0.668 0.618 0.633 0.636 0.587 0.664 0.564 0.627 0.672 0.636 0.667 0.728 0.852
${}M~{}\downarrow$ 0.107 0.084 0.102 0.097 0.109 0.085 0.099 0.096 0.092 0.105 0.087 0.111 0.095 0.083 0.092 0.085 0.070 0.035
$S_{\alpha}\uparrow$ 0.755 0.786 0.737 0.755 0.735 0.788 0.753 0.761 0.763 0.729 0.784 0.740 0.759 0.777 0.759 0.794 0.823 0.884
$E_{\phi}^{m}\uparrow$ 0.796 0.836 0.804 0.820 0.786 0.833 0.823 0.829 0.828 0.813 0.840 0.781 0.826 0.856 0.834 0.832 0.858 0.938
${\rm HCE}_{\gamma}\downarrow$ 474 480 501 542 574 490 593 567 586 516 555 621 512 600 556 451 340 480
DIS-TE3 $maxF_{\beta}\uparrow$ 0.748 0.785 0.726 0.743 0.699 0.798 0.730 0.751 0.747 0.717 0.784 0.710 0.752 0.772 0.745 0.792 0.830 0.885
$F^{w}_{\beta}\uparrow$ 0.644 0.696 0.620 0.656 0.590 0.707 0.641 0.664 0.657 0.623 0.700 0.595 0.664 0.702 0.662 0.701 0.758 0.862
${}M~{}\downarrow$ 0.098 0.083 0.103 0.092 0.109 0.079 0.096 0.092 0.092 0.102 0.080 0.109 0.091 0.078 0.090 0.079 0.064 0.035
$S_{\alpha}\uparrow$ 0.780 0.798 0.747 0.773 0.748 0.809 0.766 0.777 0.774 0.749 0.805 0.757 0.780 0.794 0.771 0.811 0.836 0.883
$E_{\phi}^{m}\uparrow$ 0.827 0.856 0.815 0.848 0.801 0.858 0.849 0.854 0.843 0.833 0.869 0.801 0.852 0.880 0.855 0.857 0.883 0.951
${\rm HCE}_{\gamma}\downarrow$ 883 948 972 1059 1058 965 1096 1082 $1$ 999 1049 1146 1001 1136 1081 887 687 973
DIS-TE4 $maxF_{\beta}\uparrow$ 0.759 0.780 0.729 0.721 0.670 0.795 0.699 0.731 0.725 0.715 0.772 0.710 0.749 0.736 0.731 0.782 0.827 0.848
$F^{w}_{\beta}\uparrow$ 0.659 0.693 0.625 0.633 0.559 0.705 0.616 0.647 0.630 0.621 0.687 0.598 0.663 0.664 0.652 0.693 0.753 0.829
${}M~{}\downarrow$ 0.102 0.091 0.109 0.107 0.127 0.087 0.113 0.107 0.107 0.111 0.092 0.114 0.099 0.098 0.102 0.091 0.072 0.049
$S_{\alpha}\uparrow$ 0.784 0.794 0.743 0.752 0.723 0.807 0.744 0.763 0.758 0.744 0.792 0.755 0.776 0.770 0.762 0.802 0.830 0.854
$E_{\phi}^{m}\uparrow$ 0.821 0.848 0.803 0.825 0.767 0.847 0.824 0.838 0.815 0.820 0.854 0.788 0.837 0.848 0.841 0.842 0.870 0.938
${\rm HCE}_{\gamma}\downarrow$ 3218 3601 3654 3760 3678 3653 3683 3803 3806 3709 3864 3999 3690 3817 3819 3331 2888 3799
Overall DIS-TE (1-4) $maxF_{\beta}\uparrow$ 0.708 0.752 0.694 0.704 0.660 0.761 0.693 0.712 0.710 0.678 0.743 0.674 0.711 0.729 0.710 0.757 0.799 0.863
$F^{w}_{\beta}\uparrow$ 0.603 0.663 0.590 0.614 0.554 0.670 0.608 0.624 0.620 0.584 0.658 0.558 0.622 0.658 0.628 0.665 0.726 0.839
${}M~{}\downarrow$ 0.103 0.086 0.103 0.098 0.112 0.083 0.101 0.097 0.095 0.105 0.087 0.110 0.095 0.085 0.094 0.084 0.070 0.039
$S_{\alpha}\uparrow$ 0.759 0.783 0.732 0.750 0.728 0.791 0.747 0.756 0.755 0.729 0.781 0.737 0.758 0.770 0.754 0.792 0.819 0.872
$E_{\phi}^{m}\uparrow$ 0.798 0.835 0.797 0.819 0.776 0.835 0.822 0.827 0.819 0.810 0.840 0.778 0.825 0.850 0.832 0.834 0.858 0.936
${\rm HCE}_{\gamma}\downarrow$ 1202 1313 1339 1401 1395 1333 1411 1427 1442 1365 1432 1513 1359 1457 1426 1218 1016 1364
@@
More Qualitative Evaluation. As shown in Fig. 13, our method yields more refined segmentation results, providing cleaner foreground masks. It also produces precise outputs for intricate lines that are challenging for previous approaches to handle.
@@
Cross Dataset Evaluation. To test the generalization ability of our model, we randomly select some images from other datasets(Agustsson & Timofte, 2017; Lin et al., 2014; Shao et al., 2019) for experiments. As shown in Fig. 10, compared to IS-Net and IS-Net-General-UseQin et al. (2022), our approach exhibits finer segmentation quality across diverse images, providing cleaner foreground masks. It is noteworthy that IS-Net-General-Use is fine-tuned on extra datasets to enhance generalization, which indicates that our method has a stronger generalization ability.

### v4

@@ Appendix E More Visualization Analysis
Quantitative Comparisons of Ablations. Visualization of the ablation study experiments in the main paper §2 is shown in Fig. 7. The estimated depth detail remains comparable with a customized “DPT head” (Ranftl et al., 2021). Models trained with lower-quality data like “Taskonomy + Cityscapes” or without the pre-trained VAE decoder parameters suffer from a decline in the ability to predict details. We attribute them to the low-quality annotation and the oversized decoder, respectively.
@@
Forward Diffusion Process. More detailed visualization of different noise forms and proportions in the forward diffusion process is shown in Fig. 8. With larger ($\beta_{start}$, $\beta_{end}$) values of the DDIM scheduler, the proportion of noise will be much higher during the forward diffusion process. When ($\beta_{start}$, $\beta_{end}$) reaches (1.0, 1.0), the noisy latent will be pure noise latent, which is Gaussian noise and RGB latent for (a) and (b), respectively.
Quantitative Comparisons of Generalization. We compare the generalization performance of models trained on synthetic (50K Hypersim (Roberts et al., 2021) + 40K Virtual KITTI (Cabon et al., 2020)) and real data (50K Taskonomy (Zamir et al., 2018) + 40K Cityscapes (Cordts et al., 2016)) for out-of-distribution scenarios. As illustrated in Fig. 9, the robustness of GenPercept trained on real data is comparable to that trained on synthetic data in challenging scenes, such as underwater environments, non-realistic renderings, and evening settings. Notably, models trained on synthetic data demonstrate superior accuracy in estimating transparent objects and capturing geometric details, owing to the high-quality and densely labeled synthetic ground truth.
Figure 9: Quantitative comparisons of generalization for models trained on synthetic and real data.
Appendix F An Extra Attempt on Human Pose Estimation
@@
Appendix G Implementation Details
Implementation Details of Exploration Experiments in Section 3 of the Main Paper. Unless specified otherwise, we follow Marigold’s training setting and train for 30000 iterations to estimate the affine-invariant monocular depth. The training dataset contains 50K Hypersim (Roberts et al., 2021) images and 40K Virtual KITTI (Cabon et al., 2020) images, and these images are sampled with a sampling rate of 90% for Hypersim and 10% for Virtual KITTI. We freeze the VAE AutoEncoder and fine-tune the U-Net of Stable Diffusion v2.1 to estimate the ground-truth label latent, with a resolution of (480, 640), a batch size of 32, and a learning rate of 3e-5. The multi-resolution noise (Ke et al., 2024) is employed for Gaussian noise and not used for RGB noise (Lee et al., 2024) by default. For inference, the ensemble size and denoising steps are set to 1 and 10, respectively. Evaluation results of absolute relative error (AbsRel) and $\delta_{1}$ are reported on five unseen monocular depth datasets, including KITTI (Geiger et al., 2013), NYU (Silberman et al., 2012), ScanNet (Dai et al., 2017), DIODE (Vasiljevic et al., 2019), and ETH3D (Schops et al., 2017).
Appendix H More Quantitative and Qualitive Experiments
H.1 Monocular Depth Estimation
@@
H.2 Surface Normal Estimation
@@
H.3 Dichotomous Image Segmentation
More Quantitative Evaluation. To conduct a comprehensive evaluation, we compare our approach with numerous previous models including models for medical image segmentation (Ronneberger et al., 2015), salient object detection (Qin et al., 2019; Zhao et al., 2020; Wei et al., 2020; Chen et al., 2020; Qin et al., 2020), camouflaged object detection (Fan et al., 2021a; Mei et al., 2021), semantic segmentation (Zhao et al., 2017; Chen et al., 2018a; Wang et al., 2020; Yu et al., 2018; Zhao et al., 2018; Howard et al., 2019; Fan et al., 2021c; Nirkin et al., 2021) and models like IS-Net (Qin et al., 2022) and MVANet (Yu et al., 2024) specifically trained for DIS.
As shown in Table 14, our proposed model significantly outperforms methods like IS-Net across most evaluation metrics on this challenging dataset. Compared with MVANet, which inferences with a resolution of $1024\times 1024$, our GenPercept achieves slightly lower performance, but the results remains competitive and can highlight the effectiveness of our approach for DIS.
@@ H.3 Dichotomous Image Segmentation
Ours (1024px)
DIS-VD $maxF_{\beta}\uparrow$ 0.692 0.731 0.678 0.685 0.648 0.748 0.665 0.691 0.691 0.660 0.726 0.662 0.697 0.714 0.696 0.734 0.791 0.904 0.857 0.877
$F^{w}_{\beta}\uparrow$ 0.586 0.641 0.574 0.595 0.542 0.656 0.584 0.604 0.603 0.568 0.641 0.548 0.609 0.642 0.613 0.640 0.717 0.861 0.835 0.859
${}M~{}\downarrow$ 0.113 0.094 0.110 0.107 0.118 0.090 0.110 0.106 0.102 0.114 0.095 0.116 0.102 0.092 0.103 0.096 0.074 0.035 0.040 0.035
$S_{\alpha}\uparrow$ 0.745 0.768 0.723 0.733 0.718 0.781 0.727 0.740 0.744 0.716 0.767 0.728 0.747 0.758 0.740 0.773 0.813 0.909 0.870 0.887
$E_{\phi}^{m}\uparrow$ 0.785 0.816 0.783 0.800 0.765 0.823 0.798 0.811 0.802 0.796 0.824 0.767 0.811 0.841 0.817 0.814 0.856 0.937 0.934 0.941
${\rm HCE}_{\gamma}\downarrow$ 1337 1402 1493 1567 1555 1413 1568 1606 1588 1520 1560 1660 1503 1625 1598 1324 1116 878 1511 1262
DIS-TE1 $maxF_{\beta}\uparrow$ 0.625 0.688 0.620 0.640 0.598 0.694 0.644 0.646 0.645 0.601 0.668 0.595 0.631 0.669 0.648 0.695 0.740 0.893 0.841 0.850
$F^{w}_{\beta}\uparrow$ 0.514 0.595 0.517 0.549 0.495 0.601 0.558 0.552 0.557 0.506 0.579 0.474 0.535 0.595 0.562 0.597 0.662 0.823 0.814 0.827
${}M~{}\downarrow$ 0.106 0.084 0.099 0.095 0.103 0.083 0.094 0.094 0.089 0.102 0.088 0.108 0.095 0.083 0.090 0.082 0.074 0.037 0.038 0.036
$S_{\alpha}\uparrow$ 0.716 0.754 0.701 0.721 0.705 0.760 0.727 0.722 0.725 0.694 0.742 0.695 0.716 0.740 0.723 0.761 0.787 0.879 0.868 0.878
$E_{\phi}^{m}\uparrow$ 0.750 0.801 0.766 0.783 0.750 0.801 0.791 0.786 0.791 0.772 0.797 0.741 0.784 0.818 0.798 0.803 0.820 0.911 0.918 0.919
${\rm HCE}_{\gamma}\downarrow$ 233 220 230 244 271 224 274 253 267 234 262 288 234 274 249 205 149 103 204 165
DIS-TE2 $maxF_{\beta}\uparrow$ 0.703 0.755 0.702 0.712 0.673 0.756 0.700 0.720 0.724 0.681 0.747 0.680 0.716 0.743 0.720 0.759 0.799 0.925 0.876 0.880
$F^{w}_{\beta}\uparrow$ 0.597 0.668 0.598 0.620 0.570 0.668 0.618 0.633 0.636 0.587 0.664 0.564 0.627 0.672 0.636 0.667 0.728 0.874 0.852 0.859
${}M~{}\downarrow$ 0.107 0.084 0.102 0.097 0.109 0.085 0.099 0.096 0.092 0.105 0.087 0.111 0.095 0.083 0.092 0.085 0.070 0.03 0.035 0.034
$S_{\alpha}\uparrow$ 0.755 0.786 0.737 0.755 0.735 0.788 0.753 0.761 0.763 0.729 0.784 0.740 0.759 0.777 0.759 0.794 0.823 0.915 0.884 0.892
$E_{\phi}^{m}\uparrow$ 0.796 0.836 0.804 0.820 0.786 0.833 0.823 0.829 0.828 0.813 0.840 0.781 0.826 0.856 0.834 0.832 0.858 0.944 0.938 0.938
${\rm HCE}_{\gamma}\downarrow$ 474 480 501 542 574 490 593 567 586 516 555 621 512 600 556 451 340 246 480 410
DIS-TE3 $maxF_{\beta}\uparrow$ 0.748 0.785 0.726 0.743 0.699 0.798 0.730 0.751 0.747 0.717 0.784 0.710 0.752 0.772 0.745 0.792 0.830 0.936 0.885 0.898
$F^{w}_{\beta}\uparrow$ 0.644 0.696 0.620 0.656 0.590 0.707 0.641 0.664 0.657 0.623 0.700 0.595 0.664 0.702 0.662 0.701 0.758 0.89 0.862 0.879
${}M~{}\downarrow$ 0.098 0.083 0.103 0.092 0.109 0.079 0.096 0.092 0.092 0.102 0.080 0.109 0.091 0.078 0.090 0.079 0.064 0.031 0.035 0.032
$S_{\alpha}\uparrow$ 0.780 0.798 0.747 0.773 0.748 0.809 0.766 0.777 0.774 0.749 0.805 0.757 0.780 0.794 0.771 0.811 0.836 0.92 0.883 0.896
$E_{\phi}^{m}\uparrow$ 0.827 0.856 0.815 0.848 0.801 0.858 0.849 0.854 0.843 0.833 0.869 0.801 0.852 0.880 0.855 0.857 0.883 0.954 0.951 0.954
${\rm HCE}_{\gamma}\downarrow$ 883 948 972 1059 1058 965 1096 1082 $1$ 999 1049 1146 1001 1136 1081 887 687 512 973 809
DIS-TE4 $maxF_{\beta}\uparrow$ 0.759 0.780 0.729 0.721 0.670 0.795 0.699 0.731 0.725 0.715 0.772 0.710 0.749 0.736 0.731 0.782 0.827 0.911 0.848 0.874
$F^{w}_{\beta}\uparrow$ 0.659 0.693 0.625 0.633 0.559 0.705 0.616 0.647 0.630 0.621 0.687 0.598 0.663 0.664 0.652 0.693 0.753 0.857 0.829 0.858
${}M~{}\downarrow$ 0.102 0.091 0.109 0.107 0.127 0.087 0.113 0.107 0.107 0.111 0.092 0.114 0.099 0.098 0.102 0.091 0.072 0.041 0.049 0.041
$S_{\alpha}\uparrow$ 0.784 0.794 0.743 0.752 0.723 0.807 0.744 0.763 0.758 0.744 0.792 0.755 0.776 0.770 0.762 0.802 0.830 0.903 0.854 0.874
$E_{\phi}^{m}\uparrow$ 0.821 0.848 0.803 0.825 0.767 0.847 0.824 0.838 0.815 0.820 0.854 0.788 0.837 0.848 0.841 0.842 0.870 0.944 0.938 0.947
${\rm HCE}_{\gamma}\downarrow$ 3218 3601 3654 3760 3678 3653 3683 3803 3806 3709 3864 3999 3690 3817 3819 3331 2888 2301 3799 3321
Overall DIS-TE (1-4) $maxF_{\beta}\uparrow$ 0.708 0.752 0.694 0.704 0.660 0.761 0.693 0.712 0.710 0.678 0.743 0.674 0.711 0.729 0.710 0.757 0.799 0.916 0.863 0.875
$F^{w}_{\beta}\uparrow$ 0.603 0.663 0.590 0.614 0.554 0.670 0.608 0.624 0.620 0.584 0.658 0.558 0.622 0.658 0.628 0.665 0.726 0.855 0.839 0.856
${}M~{}\downarrow$ 0.103 0.086 0.103 0.098 0.112 0.083 0.101 0.097 0.095 0.105 0.087 0.110 0.095 0.085 0.094 0.084 0.070 0.035 0.039 0.036
$S_{\alpha}\uparrow$ 0.759 0.783 0.732 0.750 0.728 0.791 0.747 0.756 0.755 0.729 0.781 0.737 0.758 0.770 0.754 0.792 0.819 0.905 0.872 0.885
$E_{\phi}^{m}\uparrow$ 0.798 0.835 0.797 0.819 0.776 0.835 0.822 0.827 0.819 0.810 0.840 0.778 0.825 0.850 0.832 0.834 0.858 0.938 0.936 0.939
${\rm HCE}_{\gamma}\downarrow$ 1202 1313 1339 1401 1395 1333 1411 1427 1442 1365 1432 1513 1359 1457 1426 1218 1016 790 1364 1176
@@
More Qualitative Evaluation. As shown in Fig. 14, our method yields refined segmentation results, providing cleaner foreground masks. It also can produce precise outputs for intricate lines.
@@
Cross Dataset Evaluation. To test the generalization ability of our model, we randomly select some images from other datasets(Agustsson & Timofte, 2017; Lin et al., 2014; Shao et al., 2019) and in-the-wild images for experiments. As shown in  Fig. 13, compared to IS-Net and IS-Net-General-Use(Qin et al., 2022), our approach exhibits finer segmentation quality across diverse images, providing cleaner foreground masks. Compared to MVANet (Yu et al., 2024), GenPercept exhibits enhanced robustness when applied to in-the-wild images. This improvement can be attributed to its large-scale pre-training on the LAION dataset and the extensive parameterization of the diffusion model. In contrast, the backbone of MVANet is pre-trained on the ImageNet dataset (Deng et al., 2009), which may limit its performance in more diverse environments.
It is noteworthy that IS-Net-General-Use is fine-tuned on extra datasets to enhance generalization, which indicates that our method has a stronger generalization ability.
H.4 Image Matting
More Qualitative Evaluation. In Fig. 15, we showcase more qualitative results for the image matting task. It is worth noting that our model works well in various resolutions, light environments, human poses, and human orientations. More importantly, our GenPercept model trained on human matting images shows much more robustness to other objects compared to existing P3M10K (Li et al., 2021) SOTA method ViTAE-S (Ma et al., 2023), as illustrated in Fig. 16. It shows that the ViTAE-S overfits the human matting task, while GenPercept preserves the generalization ability. Besides, we also train GenPercept on a more general image matting task on the Composition-1K (Xu et al., 2017) dataset. As shown in Fig. 17, GenPercept shows robustness on more types of objects such as semi-transparent objects, hollow objects, etc.
H.5 Image Segmentation
H.6 Human Pose Estimation
More Qualitative Evaluation. In Fig. 19, we showcase more qualitative results for the human pose estimation. We conduct experiments on MHP dataset (Li et al., 2017), and we use mmpose (Contributors, 2020) to render the human pose following the setting of (Bai et al., 2023).
Figure 11: More qualitative results for monocular depth estimation.
Figure 12: More qualitative comparisons of surface normal estimation. Our GenPercept can achieve more detailed results, even compared to the competitive CVPR2024 method DSINE (Bae & Davison, 2024b). Note that these two visualization coordinate systems are slightly different.